OpenStudy (anonymous):

how can i place the following numbers on a pie chart? Chocholate = 45.32% Milk = 23.76% Eggs = 15.46% Cherries = 15.46% I know its very easy to plot some stuff on a pie chart...but these are harder as they have stuff like .46%....

OpenStudy (mathstudent55):

Each amount you need to place on the pie chart has to represent a number of degrees. Since a complete circle has 360 degrees, take all those percentages of 360 to find how many degrees each one is. For example: Chocolate = 45.32% 45.32% of 360 degrees = 0.4532 * 360 degrees = 163 degrees Cherries = 15.46% 15.46% of 360 degrees = 0.1546 * 360 degrees = 56 degrees Do the same for milk and eggs to get their numbers of degrees.

OpenStudy (anonymous):

thank you very much,i've managed to deal it :D

OpenStudy (mathstudent55):

Great. You're welcome.